# No Minima, No Collisions: Combining Modulation and Control Barrier Function Strategies for Feasible Dynamical Collision Avoidance

### v1

@@ IV-B Qualitative Comparison: Obstacle Geometry and Trapping Region
Mod-DS approaches’ performances differ given their choices of $\lambda(x,t)$, $\lambda_{e}(x,t)$ and $d(x,t)$. Likewise, the parameterization of $\alpha$ functions CBF-QP affects the obstacle-avoiding behavior and speed profiles of the safe controller. In Figure 2 we present stream plots depicting the safe DS $\dot{x}$ modified by CBP-QP and Mod-DS approaches respectively, given nominal DS in (23). Note that $\dot{x}=u$ in fully actuated systems (10).
@@ IV-B1 Preservation of Nominal DS
By comparing the color, which indicates the magnitude of modified to nominal control output ratios, and curvature, which represents directions of the velocities $u$ distorted by safe controllers at each state $x$, of the streamlines in Figure 2, it can be qualitatively concluded that CBF-QP with $\alpha(h)=h$ and Mod-DS’s ability in preserving the nominal DS is weaker than CBF-QP with $\alpha(h)=5h$, because streamlines or trajectories generated by CBF-QP with $\alpha(h)=5h$ are straighter, i.e. more similar to that in nominal linear DS (23), and have larger yellow region. The color yellow means the magnitude of the modified controller $||u||_{2}$ is the same as that of the nominal ones $||u_{\text{nom}}|||_{2}$ at that state $x$. In CBF-QP, the larger $\alpha(h)$ is, given input $h=h(x,t)$, the smaller the lower bound of $\dot{h}(x,t)$ is, which results in a larger feasible set for $u$. Mod-DS approaches are weaker in preserving the nominal DS than CBF-QP, which can be qualitatively observed from the coverage of red regions in Figure 2. Mod-DS modified controller often becomes larger in magnitude than the nominal controller because parameter $\lambda_{e}$ (Table I) would boost $\dot{x}$ in the directions tangent to the obstacles. Additionally, viewing from the streamlines’ curvatures in row 7, on-manifold Mod-DS drastically alters and even reverses the directions of $u_{\text{nom}}$.
@@ IV-B2 Obstacle Geometries
While CBF-QP and Mod-DS differ in their ability to preserve the nominal DS, they are similarly capable of handling convex obstacles. The generated trajectories form no obvious trapping regions, where the agent could be stuck, except an undesirable equilibrium at the boundary of the obstacle where $u_{\text{nom}}$ and $\nabla_{x}h(x,t)$ are collinear. Further analysis of this undesirable equilibrium can be found in subsection IV-C. Plots 1(c)-5(c) and 1(e)-5(e) suggest the formation of a region of attraction when using CBF-QP to avoid star-shaped and non-star-shaped concave obstacles. Agents in the trapping region are led by one side of the obstacle boundary to move left and the other side to move right, which leads to convergence at an undesirable equilibrium point on the obstacle boundary. While the normal Mod-DS on row 5 forms the same trapping region as CBF-QP approaches, reference Mod-DS navigates star-shaped obstacles without forming any region of attraction and on-manifold Mod-DS is trapping-region-free for all obstacle geometries. Thus, the conclusion can be qualitatively drawn that reference and on-manifold Mod-DS provide more reasonable solutions facing concave obstacles.
@@ IV-C Quantitative Behavior Comparison
As reactive safe control methods, both Mod-DS and CBF-QP have quick runtime, making them useful in highly dynamic environments. Mod-DS approaches’ strength is their high flexibility and adaptation to all obstacle geometries using different eigenvector and eigenvalue combinations. That ability enables Mod-DS variants like on-manifold Mod-DS to realize local minimum free obstacle avoidance, proven by its 100% success rate in converging to the target in Table II. Additionally, on-manifold Mod-DS has the highest obstacle clearance $d_{\text{obs}}$ among all, indicating higher safety guarantees with less chance of collisions. on-manifold Mod-DS’s drawback is slower runtime (though still fast enough to be reactive), the high demand for jerk $j$ when reshaping the flow, longer trajectories $l$, and larger straight-line deviation $\eta$. On the contrary, CBF-QP methods are capable of generating smooth and efficient trajectories with low $j$ and low $v_{\text{near}}$ and low length ratios $l/l_{\text{nom}}$, while their weaknesses are small target reaching rate, which reveals the existence of numerous local minima in the CBF-QP modified trajectories. Normal Mod-DS’s performance metrics are similar to that of the CBF-QP, except that the method’s popular choice of eigenvalues $\lambda_{e}>1$ (see Table I) results in higher robot velocities in the direction tangent to the boundary set $\partial C$ and a larger chance for the robot to move away from local minima on concave obstacles’ boundaries. Therefore, normal Mod-DS metrics have higher near-obstacle velocities $v_{\text{near}}$ and slightly higher convergence rate than CBF-QPs. Lastly, Reference Mod-DS’s behaviors are similar to normal Mod-DS in length ratio $l/l_{\text{nom}}$, obstacle clearance $d_{\text{obs}}$ and near-obstacle velocity $v_{\text{near}}$. But the replacement of normal vector $n$ with reference vector $r$ gives reference Mod-DS a better convergence rate for non-convex obstacle avoidance.
@@ V Theoretical Connections behind Mod-DS and CBF-QP in Fully-Actuated Systems
Qualitative and quantitative analysis of Mod-DS and CBF-QP approaches for fully actuated systems reveals some common features shared by the two seemingly distinctive approaches: given the same input state $x$, static obstacle environment $h(x)$ and nominal controller $u_{\text{nom}}$, normal Mod-DS shares the same local minimum set with CBF-QP. In this section, we demonstrate with mathematical derivations the theoretical connections responsible for similar behaviors in Mod-DS and CBF-QP performances. We prove that normal Mod-DS is equivalent to CBF-QP under certain choices of $\lambda$ and $\lambda_{e}$ in (16) and (18). We also show that solutions from reference Mod-DS are guaranteed to satisfy CBF conditions in (7) and the relationship between reference Mod-DS and CBF-QP can be mathematically quantified to convert one from another. This conversion formula is important because it shows the key closed-form components in reference Mod-DS solutions that make possible navigation in star-shaped obstacle environments with few local minima. In section VII, through inverse engineering from the conversion formula, modifications of CBF-QP that enable concave unsafe set navigation with fewer and even no local minima are proposed. Note that in this section, our theoretical analysis focuses on static obstacle environment settings. Hence, $h(x)$, $n(x)$, and $r(x)$ are used instead of $h(x,t)$, $n(x,t)$, and $r(x,t)$ to represent boundary functions, normal vectors and reference vectors.
@@ V-C Reference Mod-DS and CBF-QP Similarities
In subsection IV-B, the performance of the reference Mod-DS was demonstrated to be superior to that of the CBF-QP. Still, the fact that reference Mod-DS shared the same undesirable local minimum with CBF-QPs for circular obstacle avoidance in Figure 3 suggests the potential for hidden connections between the 2 seemingly distinctive approaches. Quantifying the similarities and differences between reference Mod-DS and CBF-QP is key to understanding how undesirable local minima can be reduced or even eliminated to realize global convergence. We begin our discussion here by proving $u_{\text{mod}}$ from reference Mod-DS satisfies the CBF constraints at all times and then quantify what changes need to be introduced before reference Mod-DS and CBF-QP can be converted to each other. The theoretical connections discussed here will contribute to the design of a new safe control approach that incorporates the advantage of reference Mod-DS and CBF-QP in section VII.
@@
Given fully-actuated nominal controller $u_{\text{nom}}$ in (10) and any $\lambda$ and $\lambda_{e}$ satisfying (18), there exists an extended class $K_{\infty}$ function such that the Mod-DS safe controller $u_{\text{mod}}$ defined in Eq. (14), (16), and (19) meets the CBF conditions defined in Eq. (11).
Proof: Since we assume that the unsafe sets are translating and rotating rather than expanding, there must exist a point $x_{o}$ somewhere such that the motion of an unsafe set can be characterized about. Thus, $\frac{\partial h(x,t)}{\partial t}$ can be rewritten as in (30), where $\dot{x}_{o}$ can be computed as in (15).
$\frac{\partial h(x,t)}{\partial t}=\frac{\partial h(x-x_{o}(t))}{\partial t}$ $=\nabla_{x_{o}}h(x-x_{o}(t))^{\top}\dot{x}_{o}$ (30)
$=-\nabla_{x}h(x-x_{0}(t))^{\top}\dot{x}_{o}$
$=-\nabla_{x}h(x,t)^{\top}\dot{x}_{o}.$
Plug (30) into the CBF constraint in (11), CBF conditions in fully actuated systems are equivalent to the following. Denote $h(x,t)=h$ and $\nabla_{x}h=\nabla_{x}h(x,t)$.
$\frac{\nabla_{x}h^{\top}}{||\nabla_{x}h||_{2}}u\geq\frac{\nabla_{x}h^{\top}}{||\nabla_{x}h||_{2}}\dot{x}_{o}-\frac{\alpha(h)}{{||\nabla_{x}h||_{2}}}$ (31)
Following the same notation as in Theorem V.4, $\forall x\in\mathbb{R}^{d}$, denote $H=[e_{1}(x,t),...e_{d-1}(x,t)]\in\mathbb{R}^{d\times(d-1)}$ to be the hyperplane plane perpendicular to the gradient of the boundary function $\nabla_{x}h$. Denote $r=r(x,t)$, $e_{i}=e_{i}(x,t),i\in\{1,2,...,d-1\}$ and the basis matrix of reference Mod-DS $E_{r}=[r(x,t),H(x,t)]$, as defined in (16) and (19). Note that $||r(x,t)||_{2}=1$ by definition. Note that since $Er(x,t)$ is an orthonormal basis, $E_{r}^{-1}\neq E_{r}^{\top}$. Instead,
@@
Therefore, the decomposition of $u_{\text{mod}}$ onto $r(x,t)$ by basis matrix $E_{r}$ can be computed from Eq. (14) and (16) as
$\quad\;u_{\text{mod}}^{r}$ (33)
$=\lambda E_{r}^{-1}[1,:](u_{\text{nom}}-\dot{x}_{o})+E_{r}^{-1}[1,:]\dot{x}_{o}$
$=\frac{\lambda}{\nabla_{x}h^{\top}r}\nabla_{x}h^{\top}u_{\text{nom}}+\frac{1-\lambda}{\nabla_{x}h^{\top}r}\nabla_{x}h^{\top}\dot{x}_{o}$
$=\frac{\nabla_{x}h^{\top}}{\nabla_{x}h^{\top}r}[\lambda u_{\text{nom}}+(1-\lambda)\dot{x}_{o}].$
The decomposition of $u_{\text{mod}}$ onto $r(x,t)$ can be further broken down in direction $n(x,t)$ parallel to $\nabla_{x}h$ and direction $e=e(x,t)$ on the tangent hyperplane $H(x,t)$. The projection of $u_{\text{mod}}$ onto $n(x,t)$ is equivalent to the projection of $u_{\text{mod}}$ first onto $r(x,t)$ and then onto $n$, i.e. the projection of $u_{\text{mod}}^{r}r(x,t)$ onto $n(x,t)$, because the vector $e(x,t)$ is orthogonal to $n(x,t)$, i.e. $\langle n(x,t),e(x,t)\rangle=0$.
$\quad\;u_{\text{mod}}^{n}=n^{\top}(u_{\text{mod}}^{r}r+u_{\text{mod}}^{H}e)=n^{\top}u_{\text{mod}}^{r}r$ (34)
$=n^{\top}r\{\frac{\nabla_{x}h^{\top}}{\nabla_{x}h^{\top}r}[\lambda u_{\text{nom}}+(1-\lambda)\dot{x}_{o}]\}$
From the relationship between vector $n(x,t)$ and $r(x,t)$ depicted in 5(b), it can be deduced that
$\begin{split}\langle\hat{n},\hat{r}\rangle=\frac{\nabla_{x}h^{\top}}{||\nabla_{x}h||_{2}}r=\cos\theta.\end{split}$ (35)
Since $-\pi/2\leq\theta\leq\pi/2$ under all conditions because the reference direction $r(x,t)$ and the normal direction $n(x,t)$ must both point away from the obstacle, $\cos\theta\geq 0$ always holds. Therefore, plugging (35) into (34),
$\quad\;u_{\text{mod}}^{n}$ $=\cos{\theta}\{\frac{\nabla_{x}h^{\top}}{\nabla_{x}h^{\top}r}[\lambda u_{\text{nom}}+(1-\lambda)\dot{x}_{o}]\}$ (36)
$=\frac{\nabla_{x}h^{\top}}{||\nabla_{x}h||_{2}}[\lambda u_{\text{nom}}+(1-\lambda)\dot{x}_{o}].$
Now we are left with showing the modulated value in (34) satisfies the CBF condition in (11). WTS: $\forall$ possible $\nabla_{x}h(x,t),h(x,t),u_{\text{nom}},\dot{x}_{o}$ and $\lambda$ combinations, there exists a $K_{\infty}$ function $\alpha$ such that the following inequility, acquired by plugging $\frac{\nabla_{x}h^{\top}}{||\nabla_{x}h||_{2}}u=u_{\text{mod}}^{n}$ from (36) into (31) and simplify, always holds.
$\alpha(h)\geq-\lambda\nabla_{x}h^{\top}u_{\text{nom}}+\lambda\nabla_{x}h^{\top}\dot{x}_{o}$ (37)
Since the norms of $\nabla_{x}h(x,t),h(x,t),u_{\text{nom}},\dot{x}_{o}$ and $\lambda$ values are all well-bounded in real-life robot applications, find $\alpha$ that serve as an upper bound for domain $\{x\in\mathbb{R}^{d}:h(x,t)>0\}$ is trivial. When $h(x,t)=0$, $\lambda=0$ based on (18), the left and right sides of the inequality become 0 and satisfy the property of $K_{\infty}$ functions. In conclusion, given a bounded fully-actuated nominal controller $u_{\text{nom}}$ in (10) and any feasible $\lambda$ and $\lambda_{e}$, there exists an extended class $K_{\infty}$ function such that the Mod-DS safety controller $u_{\text{mod}}$ meets the CBF conditions.

### v2

@@ IV-B Qualitative Comparison: Obstacle Geometry and Trapping Region
Mod-DS approaches’ performances differ given their choices of $\lambda(x,x_{o})$, $\lambda_{e}(x,x_{o})$ and $d(x,x_{o})$. Likewise, the parameterization of $\alpha$ functions CBF-QP affects the obstacle-avoiding behavior and speed profiles of the safe controller. In Figure 2 we present stream plots depicting the safe DS $\dot{x}$ modified by CBP-QP and Mod-DS approaches respectively, given nominal DS in (30). Note that $\dot{x}=u$ in fully actuated systems (17).
@@ IV-B1 Preservation of Nominal DS
By comparing the color, which indicates the magnitude of modified to nominal control output ratios, and curvature, which represents directions of the velocities $u$ distorted by safe controllers at each state $x$, of the streamlines in Figure 2, it can be qualitatively concluded that CBF-QP with $\alpha(h)=h$ and Mod-DS’s ability in preserving the nominal DS is weaker than CBF-QP with $\alpha(h)=5h$, because streamlines or trajectories generated by CBF-QP with $\alpha(h)=5h$ are straighter, i.e. more similar to that in nominal linear DS (30), and have larger yellow region. The color yellow means the magnitude of the modified controller $||u||_{2}$ is the same as that of the nominal ones $||u_{\text{nom}}|||_{2}$ at that state $x$. In CBF-QP, the larger $\alpha(h)$ is, given input $h$, the smaller the lower bound of $\dot{h}$ is, which results in a larger feasible set for $u$. Mod-DS approaches are weaker in preserving the nominal DS than CBF-QP, which can be qualitatively observed from the coverage of red regions in Figure 2. Mod-DS modified controller often becomes larger in magnitude than the nominal controller because parameter $\lambda_{e}$ (Table I) would boost $\dot{x}$ in the directions tangent to the obstacles. Additionally, viewing from the streamlines’ curvatures in row 7, on-manifold Mod-DS drastically alters and even reverses the directions of $u_{\text{nom}}$.
@@ IV-B2 Obstacle Geometries
While CBF-QP and Mod-DS differ in their ability to preserve the nominal DS, they are similarly capable of handling convex obstacles. The generated trajectories form no obvious trapping regions, where the agent could be stuck, except an undesirable equilibrium at the boundary of the obstacle where $u_{\text{nom}}$ and $\nabla_{x}h$ are collinear. Further analysis of this undesirable equilibrium can be found in subsection IV-C. Plots 1(c)-5(c) and 1(e)-5(e) suggest the formation of a region of attraction when using CBF-QP to avoid star-shaped and non-star-shaped concave obstacles. Agents in the trapping region are led by one side of the obstacle boundary to move left and the other side to move right, which leads to convergence at an undesirable equilibrium point on the obstacle boundary. While the normal Mod-DS on row 5 forms the same trapping region as CBF-QP approaches, reference Mod-DS navigates star-shaped obstacles without forming any region of attraction and on-manifold Mod-DS is trapping-region-free for all obstacle geometries. Thus, the conclusion can be qualitatively drawn that reference and on-manifold Mod-DS provide more reasonable solutions facing concave obstacles.
@@ IV-C Quantitative Behavior Comparison
As reactive safe control methods, both Mod-DS and CBF-QP have quick runtime, making them useful in highly dynamic environments. Mod-DS approaches’ strength is their high flexibility and adaptation to all obstacle geometries using different eigenvector and eigenvalue combinations. That ability enables Mod-DS variants like on-manifold Mod-DS to realize local minimum free obstacle avoidance, proven by its 100% success rate in converging to the target in Table II. Additionally, on-manifold Mod-DS has the highest obstacle clearance $d_{\text{obs}}$ among all, indicating higher safety guarantees with less chance of collisions. on-manifold Mod-DS’s drawback is slower runtime (though still fast enough to be reactive), the high demand for jerk $j$ when reshaping the flow, longer trajectories $l$, and larger straight-line deviation $\eta$. On the contrary, CBF-QP methods are capable of generating smooth and efficient trajectories with low $j$ and low $v_{\text{near}}$ and low length ratios $l/l_{\text{nom}}$, while their weaknesses are small target reaching rate, which reveals the existence of numerous local minima in the CBF-QP modified trajectories. Normal Mod-DS’s performance metrics are similar to that of the CBF-QP, except that the method’s popular choice of eigenvalues $\lambda_{e}>1$ (see Table I) results in higher robot velocities in the direction tangent to the boundary set $\partial C_{o}$ and a larger chance for the robot to move away from local minima on concave obstacles’ boundaries. Therefore, normal Mod-DS metrics have higher near-obstacle velocities $v_{\text{near}}$ and slightly higher convergence rate than CBF-QPs. Lastly, Reference Mod-DS’s behaviors are similar to normal Mod-DS in length ratio $l/l_{\text{nom}}$, obstacle clearance $d_{\text{obs}}$ and near-obstacle velocity $v_{\text{near}}$. But the replacement of normal vector $n$ with reference vector $r$ gives reference Mod-DS a better convergence rate for non-convex obstacle avoidance.
@@ V Theoretical Connections behind Mod-DS and CBF-QP in Fully-Actuated Systems
Qualitative and quantitative analysis of Mod-DS and CBF-QP approaches for fully actuated systems reveals some common features shared by the two seemingly distinctive approaches: given the same input state $x$, static obstacle environment $h(x)$ and nominal controller $u_{\text{nom}}$, normal Mod-DS shares the same local minimum set with CBF-QP. In this section, we demonstrate with mathematical derivations the theoretical connections responsible for similar behaviors in Mod-DS and CBF-QP performances. We prove that normal Mod-DS is equivalent to CBF-QP under certain choices of $\lambda$ and $\lambda_{e}$ in (23) and (25). We also show that solutions from reference Mod-DS are guaranteed to satisfy CBF conditions in (14) and the relationship between reference Mod-DS and CBF-QP can be mathematically quantified to convert one from another. This conversion formula is important because it shows the key closed-form components in reference Mod-DS solutions that make possible navigation in star-shaped obstacle environments with few local minima. In section VII, through inverse engineering from the conversion formula, modifications of CBF-QP that enable concave unsafe set navigation with fewer and even no local minima are proposed. Note that in this section, our theoretical analysis focuses on static single-obstacle environment settings. Hence, $h(x)$, $n(x)$, and $r(x)$ are used instead of $h_{o}(x,x_{o})$, $n(x,x_{o})$, and $r(x,x_{o})$ to represent boundary functions, normal vectors and reference vectors.
@@ V-C Reference Mod-DS and CBF-QP Similarities
In subsection IV-B, the performance of the reference Mod-DS was demonstrated to be superior to that of the CBF-QP. Still, the fact that reference Mod-DS shared the same undesirable local minimum with CBF-QPs for circular obstacle avoidance in Figure 3 suggests the potential for hidden connections between the 2 seemingly distinctive approaches. Quantifying the similarities and differences between reference Mod-DS and CBF-QP is key to understanding how undesirable local minima can be reduced or even eliminated to realize global convergence. We begin our discussion here by proving $u_{\text{mod}}$ from reference Mod-DS satisfies the CBF constraints at all times in static-single-obstacle avoidance and then quantify what changes need to be introduced before reference Mod-DS and CBF-QP can be converted to each other. The theoretical connections discussed here will contribute to the design of a new safe control approach that incorporates the advantage of reference Mod-DS and CBF-QP in section VII.
@@
Given fully-actuated nominal controller $u_{\text{nom}}$ in (17) and any $\lambda$ and $\lambda_{e}$ satisfying (25), there exists an extended class $K_{\infty}$ function such that the Mod-DS safe controller $u_{\text{mod}}$ defined in Eq. (21), (23), and (26) meets the CBF conditions defined in Eq. (18) when $\dot{x}_{o}=0$.
In static-single obstacle avoidance scenarios, annotations can be simplified such that $h=h(x,x_{o})$ and $\nabla_{x}h=\nabla_{x}h(x,x_{o})$. Given $\dot{x}_{o}=\textbf{0}$ and $\bar{\dot{x}}_{o}=\textbf{0}_{d\times 1}$, CBF conditions in Eq. (18) can be rewritten as:
$\frac{\nabla_{x}h^{\top}}{||\nabla_{x}h||_{2}}u\geq-\frac{\alpha(h)}{{||\nabla_{x}h||_{2}}}.$ (37)
Following the same notation as in Theorem V.4, $\forall x\in\mathbb{R}^{d}$, denote $r=r(x,x_{o})$, $n=n(x,x_{o})$, $e_{i}=e_{i}(x,x_{o}),\forall i\in\{1,2,...,d-1\}$ and $H=[e_{1}(x,x_{o}),...e_{d-1}(x,x_{o})]\in\mathbb{R}^{d\times(d-1)}$ to be the hyperplane plane orthogonal to the gradient of the boundary function $\nabla_{x}h$. Let the basis matrix of reference Mod-DS $E_{r}=[r(x,x_{o}),H(x,x_{o})]\in\mathbb{R}^{d\times d}$, as defined in (23) and (26). Note that $||r(x,x_{o})||_{2}=1$ by definition. Since $E_{r}(x,x_{o})$ is not an orthonormal basis, $E_{r}^{-1}\neq E_{r}^{\top}$. Instead,
@@
Therefore, the decomposition of $u_{\text{mod}}$ onto $r$ by basis matrix $E_{r}$ when $\bar{\dot{x}}_{o}=0$ can be computed from Eq. (21), (23) as
$u_{\text{mod}}^{r}=\lambda E_{r}^{-1}[1,:]u_{\text{nom}}=\frac{\nabla_{x}h^{\top}}{\nabla_{x}h^{\top}r}\lambda u_{\text{nom}}.$ (39)
The decomposition of $u_{\text{mod}}$ onto $r$ can be further broken down into the direction $n$ parallel to $\nabla_{x}h$ and into the tangent hyperplane $H$ orthogonal to $\nabla_{x}h$, as depicted in 5(b). The projection of $u_{\text{mod}}^{r}$ onto $n$ is equivalent to $u_{\text{mod}}^{n}$, the projection of $u_{\text{mod}}$ onto $n$, because the vector $H$ is orthogonal to $n$, i.e. $n^{\top}H=\textbf{0}_{1\times d-1}$.
$u_{\text{mod}}^{n}$ $=n^{\top}(u_{\text{mod}}^{r}r+u_{\text{mod}}^{H}e)=u_{\text{mod}}^{r}n^{\top}r$ (40)
$=(\frac{\nabla_{x}h^{\top}}{\nabla_{x}h^{\top}r}\lambda u_{\text{nom}})(\frac{\nabla_{x}h}{||\nabla_{x}h||_{2}})^{\top}r$
$=\lambda u_{\text{nom}}\frac{\nabla_{x}h^{\top}}{||\nabla_{x}h||_{2}}$
Now we are left with showing the modulated value in (40) satisfies the CBF conditions. WTS: Given any $\nabla_{x}h,h,u_{\text{nom}}$ and $\lambda$ combinations, there exists a $K_{\infty}$ function $\alpha$ such that the following inequility, acquired by substituting $\frac{\nabla_{x}h^{\top}}{||\nabla_{x}h||_{2}}u=u_{\text{mod}}^{n}$ and $u_{\text{mod}}^{n}$ expression from (40) into (37) and simplify, always holds.
$-\lambda\nabla_{x}h^{\top}u_{\text{nom}}\leq\alpha(h)$ (41)
Since the norms of $\nabla_{x}h,u_{\text{nom}}$ and $\lambda$ values are all well-bounded by definition, find $\alpha$ that serve as an upper bound for domain $\{x\in\mathbb{R}^{d}:h(x,x_{o})>0\}$ is trivial. When $h(x,x_{o})=0$, $\lambda=0$ based on (25), the left and right sides of the inequality become 0 and satisfy the property of $K_{\infty}$ functions. In conclusion, given a bounded fully-actuated nominal controller $u_{\text{nom}}$ in (17) and any feasible $\lambda$ and $\lambda_{e}$, there exists an extended class $K_{\infty}$ function such that the Mod-DS safety controller $u_{\text{mod}}$ meets the CBF conditions.